OpenStudy (anonymous):

Ashley is digging for rocks at a geological site. She has 140 meters of rope and 4 stakes to mark off a rectangular area. Which set of dimensions will create a rectangle using all the rope Ashley has with her? 14 m × 10 m 70 m × 70 m 60 m × 10 m 55 m × 10 m

OpenStudy (anonymous):

If you find the answer I'll give medal But lest see give me a sec

OpenStudy (anonymous):

k

OpenStudy (bibby):

the formula for perimeter 2L+2W which of these fits in to that and adds up to 140?

OpenStudy (anonymous):

the first one

OpenStudy (bibby):

2(24)=48

OpenStudy (anonymous):

i still think it is the first one

OpenStudy (bibby):

you're wrong you're allowed to be wrong I won't stop you

OpenStudy (anonymous):

That's wrong

OpenStudy (anonymous):

i am sorrry i am terrible at math :(

OpenStudy (anonymous):

bibby go on

OpenStudy (anonymous):

the second?

OpenStudy (bibby):

w=70,l=70 2(70)+2(70)=140+140=280

OpenStudy (anonymous):

Your just guessing now

OpenStudy (anonymous):

let me think

OpenStudy (anonymous):

bibby I didn't get 280 I got 285

OpenStudy (bibby):

you did something wrong then lol

OpenStudy (anonymous):

oh nvm I got it ur correct

OpenStudy (anonymous):

the third one i am positive no guessing

OpenStudy (anonymous):

R you working the problem out with bibby

OpenStudy (bibby):

you're right yay you

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

Yeah you got it

OpenStudy (anonymous):

thax huys for the help i will fan thiose that help

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

lol bye

OpenStudy (bibby):

If you're coming here over a year later, the answer is C. we want the choice whose perimeter adds up to 140 C. 60x10 =>\( 60+60+10+10=140 \checkmark\)